# Non-equilibrium dynamics and atom-pair coherence in strongly interacting Bose-Fermi mixtures

### v1

@@ I INTRODUCTION
In recent years, ultracold atomic gases near magnetic Feshbach resonances have emerged as a popular platform for probing non-equilibrium physics, both due to their excellent controllability in experiments, and their amicability to accurate theoretical modelling Chin et al. (2010); Kokkelmans et al. (2002); Kokkelmans and Holland (2002); Gurarie and Radzihovsky (2007). Here, non-Gaussian correlations are essential for ergodicity Van Regemortel et al. (2018), and can arise out of the presence of non-perturbative few-body physics, including novel few-body composites bound by the medium, and the Efimov effect Levinsen et al. (2015); Fletcher et al. (2017); Klauss et al. (2017); Colussi et al. (2018); Musolino et al. (2019); Christianen et al. (2022); Musolino et al. (2022). Recently, such effects were probed via a series of experiments exploring interaction quenches of degenerate Bose gases Makotyn et al. (2014); Eigen et al. (2017, 2018), where ergodicity is required for the system to transition from the degenerate to thermal regime. Inspired by these experiments, theoretical models were formulated based on a cumulant expansion of the interparticle correlations that develop following the quench Fricke (1996); Kira and Koch (2012); Kira (2015); Colussi et al. (2020); Ahmed-Braun et al. (2022), which were finally able to connect the relaxation dynamics in these systems to the prevalence of specific non-Gaussian correlations Colussi et al. (2020); Musolino et al. (2022); van de Kraats et al. (2024).
@@
In the context of non-equilibrium and beyond-Gaussian physics, Bose-Fermi mixtures are especially interesting as the resonant vacuum two-body physics in the gas can only be captured by non-Gaussian correlations Bortolotti et al. (2005, 2008). This make Bose-Fermi mixtures distinct from identical Bose or two-component Fermi gases, where a many-body description that incorporates the two-body physics is already obtained at the level of the Gaussian Hartree-Fock-Bogoliubov approximation Blaizot and Ripka (1985). In addition, an interaction quench of a degenerate Bose-Fermi mixture necessarily induces the immediate creation of correlated Bose-Fermi pairs at nonzero center-of-mass momenta, arising from the interaction between the zero momentum Bose-Einstein condensate and the Fermi sphere. This contrasts with the unitary Bose gas, where pair excitations from the condensate can only create zero-momentum pairs, validating the Gaussian approximation at early times. Due to these theoretical difficulties, previous studies of the dynamics of Bose-Fermi mixtures rely on various simplifications, such as large population imbalances or reductions of the dimensionality of the system Wouters et al. (2003); Bortolotti et al. (2005); Olsen et al. (2009); Cumby et al. (2013); Gärttner et al. (2019); Mistakidis et al. (2019); Abdullaev et al. (2019).
@@ II.1 Hamiltonian
where the primes denote annihilation of bosons and fermions in an energetically closed spin channel, and where $\phi(q)$ is the wave function of the closed-channel molecule as a function of the relative two-body momentum $q\equiv\absolutevalue{\mathbf{q}}$ |. The molecules have kinetic energy $\varepsilon_{\mathbf{Q}}^{M}=Q^{2}/2M$, and magnetic field dependent internal energy $\nu$. A Bose-Fermi pair can associate to form a molecule, controlled by a separable interaction potential with coupling strength $g$ and form factor $\xi_{\mathbf{q}}=\Theta(\Lambda_{BF}-\absolutevalue{\mathbf{q}})$. Here $\Theta(x)$ is the Heaviside step function such that $\Lambda_{BF}$ defines an ultraviolet cutoff on the relative momentum, interpreted physically as the inverse of the range of the interspecies interaction.
@@ II.2 Cumulant expansion and equations of motion
Cumulant Notation Role
$\expectationvalue*{\hat{b}_{\mathbf{0}}}_{\mathrm{c}}$ $\sqrt{V}\psi$ Condensate density
$\expectationvalue*{\hat{b}_{\mathbf{k}}^{\dagger}\hat{b}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{B}$ Bosonic excitation density
$\expectationvalue*{\hat{f}_{\mathbf{k}}^{\dagger}\hat{f}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{F}$ Fermionic density
$\expectationvalue*{\hat{m}_{\mathbf{k}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{M}$ Molecular density
$\expectationvalue*{\hat{b}_{\mathbf{k}}\hat{b}_{-\mathbf{k}}}_{\mathrm{c}}$ $\kappa_{\mathbf{k}}$ Bosonic pairing
$\expectationvalue*{\hat{f}_{\mathbf{k}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $\chi_{\mathbf{k}}$ Atom-molecule conversion
$\expectationvalue*{\hat{b}_{\mathbf{k}-\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}^{\prime}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $M_{\mathbf{k},\mathbf{k}^{\prime}}$ Atom-molecule conversion
$\expectationvalue*{\hat{b}_{\mathbf{k}-\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}}}_{\mathrm{c}}$ $L_{\mathbf{k},\mathbf{k}^{\prime}}$ Boson-fermion scattering
Table 1: All cumulants included in our dynamical model, with a brief description of their role in the many-body dynamics. Note that by construction, a cumulant is only defined when all bosonic momenta are nonzero, except for the singlet $\expectationvalue*{\hat{b}_{\mathbf{0}}}_{\mathrm{c}}$.
@@
Next to the densities themselves, we also need to include the atom-molecule conversion processes that will act to source the density dynamics. The lowest order such process, shown in Fig. 1(a), describes the interaction between a fermion (solid black) and a boson in the BEC (wavy red), captured by the second order cumulant, or doublet, $\chi_{\mathbf{k}}$. To complete the set of doublet cumulants we define the bosonic pairing matrix $\kappa_{\mathbf{k}}$, which captures correlated pair excitations from the BEC. Note that this set excludes the fermionic and molecular pairing matrices $\expectationvalue*{\hat{f}_{\mathbf{k}}\hat{f}_{-\mathbf{k}}}_{\mathrm{c}}$ and $\expectationvalue*{\hat{m}_{\mathbf{k}}\hat{m}_{-\mathbf{k}}}_{\mathrm{c}}$, respectively, which always vanish provided that the Hamiltonian conserves the total fermionic population.
@@
In identical Bose and two-component Fermi gases, a truncation of the cumulant expansion at the level of second order correlations, or doublets, produces the Hartree-Fock-Bogoliubov respectively Blaizot and Ripka (1985), which are well studied in the literature and known to converge to the two-body Schrödinger equation in the short-range or vacuum limits Leggett (2006). In the Bose-Fermi mixture however, the same level of approximation leads to dynamical equations that do not correctly reproduce the vacuum two-body physics, pointing out the fundamental shortcomings of the doublet approximation and the inherent non-Gaussian nature of the pairing physics in the Bose-Fermi mixture Bortolotti et al. (2005, 2008). As we will show formally in Sec. III, this problem is remedied upon extending the doublet approximation with two non-Gaussian third order cumulants (triplets), defined in the last two rows of Table 1. Here, $L_{\mathbf{k},\mathbf{k}^{\prime}}$ captures the triplet contribution to the cumulant expansion of the total two-body correlator $\expectationvalue*{\hat{b}^{\dagger}\hat{f}^{\dagger}\hat{b}\hat{f}}$, alternatively interpreted as the open-channel analogue of the doublet $\chi_{\mathbf{k}}$. In turn, $M_{\mathbf{k},\mathbf{k}^{\prime}}$ captures the generalization of the scattering process in Fig. 1(a) to the case where the boson is in an excited state, as shown in Fig. 1(b). It is a necessary inclusion in the model as it ensures the conservation of the total system energy, or $\partialderivative{t}\expectationvalue*{\hat{H}}=0$.
@@
Figure 1: Two-body scattering diagrams that contribute to the density dynamics in the Bose-Fermi mixture, where red(black) wiggly lines represent (non)condensed bosons, and single(double) straight lines represent fermions(molecules). Process (a) represents scattering of fermions with the BEC, captured by $\chi_{\mathbf{Q}}$, while process (b) represents scattering of fermions with bosonic excitations, captured by $M_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{k}}$. Here $\mathbf{Q}$ and $\mathbf{k}$ are the center-of-mass and relative momenta of the Bose-Fermi pair respectively.
@@
$\bar{n}_{\mathbf{k}}^{B}=\frac{1}{V}\sum_{\mathbf{k}^{\prime}}\zeta_{\mathbf{k}-\mathbf{k}^{\prime}}^{2}n_{\mathbf{k}^{\prime}}^{B},\qquad\bar{\kappa}=\frac{1}{V}\sum_{\mathbf{k}}\zeta_{2\mathbf{k}}^{2}\kappa_{\mathbf{k}}.$ (8)
@@
$i\dot{\psi}$ $=$ $v\left(\zeta_{0}^{2}\absolutevalue*{\psi}^{2}+2\bar{n}_{0}^{B}\right)\psi+\frac{v\psi^{*}}{V}\sum_{\mathbf{k}}\zeta_{0}\zeta_{2\mathbf{k}}\kappa_{\mathbf{k}}$ (9)
@@
In turn, the dynamics of non-condensed bosons follows from,
@@
Here we have defined the Hartree-Fock Hamiltonian and pairing field, quantifying the kinetic and interaction energies of bosonic pairs in the many-body background,
@@
$\Delta_{\mathbf{k}}$ $=$ $v\zeta_{2\mathbf{k}}\left[\zeta_{0}^{2}\psi^{2}+\bar{\kappa}\right].$ (13)
Note here that in the limit of vanishing Bose-Bose scattering length, $a_{BB}=0$, we have $h_{\mathbf{k}}^{\mathrm{hf}}=\varepsilon_{\mathbf{k}}^{B}$ and $\Delta_{\mathbf{k}}=0$. The analogous equations for the fermionic and molecular densities read,
@@
Note here that the source terms in the equations of motion for the densities can be matched consistently with the scattering diagrams in Fig. 1. The cumulant $\chi_{\mathbf{k}}$ evolves as,
@@
$+$ $\frac{g}{\sqrt{V}}\sum_{\mathbf{k}^{\prime}}\xi_{\frac{m_{F}}{M}\mathbf{k}-\mathbf{k}^{\prime}}L_{\mathbf{k},\mathbf{k}^{\prime}}^{*},$
@@ III.1 Static pair energies
Before studying dynamics, we give a brief overview of relevant time-independent solutions to the embedded two-body Schrödinger equation, both in vacuum and including the many-body background. In the former case, the Hamiltonian [Eq. (1)] can be diagonalized analytically to obtain the binding energy $E_{\mathrm{vac}}$ of the dressed Feshbach dimer, resulting in the integral equation Kokkelmans et al. (2002),
@@
In a many-body system, the Hamiltonian (1) is not analytically solveable in general. As pointed out in Ref. Dukelsky et al. (2011) however, the problem may be reduced to a solveable form by neglecting coupling terms between pairs that have nonzero center-of-mass momentum $\mathbf{Q}\neq 0$, equivalent to the replacement $\hat{m}_{\mathbf{Q}}\rightarrow\hat{m}_{0}\delta_{\mathbf{Q},0}$ in the Hamiltonian. In this case the exact eigenstates of the many-body Hamiltonian that may be generated from the initial condition contain only a single interacting pair, formed from a boson in the BEC and a fermion at $\mathbf{k}=0$. The energy spectrum of this pair is obtained from the integral equation,
$E-\nu=g^{2}\xi_{\mathbf{0}}^{2}\frac{n_{0}^{B}}{E}+\frac{g^{2}}{V}\sum_{\mathbf{q}}\xi_{\mathbf{q}}^{2}\frac{1-n_{\mathbf{q}}^{F}}{E-\varepsilon_{\mathbf{q}}^{R}}.$ (20)
See App. A for more detail. The solutions $E$ of Eq. (20) are illustrated in Fig. 2. Comparing with Eq. (19) we observe that the BEC and Fermi sea have competing effects on the energy of the pair. First, the Fermi sea blocks all momentum states $k<k_{F}$, thus raising the threshold for open-channel two-body scattering to $E_{\mathbf{0}}^{\mathrm{thr}}=k_{F}^{2}/2m_{R}$, a mechanism also responsible for Cooper pairing in metals Song et al. (2010); Song and Zhou (2011). The associated enhancement of the two-body interaction makes that the Feshbach dimer no longer disappears at resonance, but rather shifts above the vacuum threshold $E=0$ at positive scattering length $(k_{F}a_{BF})^{-1}=2/\pi$. If the scattering length is tuned further to the attractive side of the resonance, the dimer moves up through the Fermi sphere towards the raised scattering threshold. Here we emphasize that this state constitutes an exact stationary state of the many-body Hamiltonian, meaning that the dimer state has infinite lifetime across the Feshbach resonance.

### v2

@@ I INTRODUCTION
In recent years, ultracold atomic gases near magnetic Feshbach resonances have emerged as a popular platform for probing non-equilibrium physics, both due to their excellent controllability in experiments, and their amicability to accurate theoretical modelling Chin et al. (2010); Kokkelmans et al. (2002); Kokkelmans and Holland (2002); Gurarie and Radzihovsky (2007). Here, non-Gaussian correlations are essential for ergodicity Van Regemortel et al. (2018), and can arise out of the presence of non-perturbative few-body physics, including novel three-body composites bound by the medium, and strongly enhanced three-body scattering by the presence of weakly bound trimer states due to the Efimov effect Levinsen et al. (2015); Fletcher et al. (2017); Klauss et al. (2017); Colussi et al. (2018); Musolino et al. (2019); Christianen et al. (2022); Musolino et al. (2022). Recently, such effects were probed via a series of experiments exploring interaction quenches of degenerate Bose gases Makotyn et al. (2014); Eigen et al. (2017, 2018), where ergodicity is required for the system to transition from the degenerate to thermal regime. Inspired by these experiments, theoretical models were formulated based on a cumulant expansion of the interparticle correlations that develop following the quench Fricke (1996); Kira and Koch (2012); Kira (2015); Colussi et al. (2020); Ahmed-Braun et al. (2022), which were finally able to connect the relaxation dynamics in these systems to the prevalence of specific non-Gaussian correlations that encode three-body scattering physics Colussi et al. (2020); Musolino et al. (2022); van de Kraats et al. (2024).
@@
In contrast to single-component Bose gases, where the vacuum two-body scattering physics is captured fully by Gaussian correlations, the equivalent physics in Bose-Fermi mixtures cannot be obtained from a Gaussian approximation Bortolotti et al. (2005, 2008). This means that, while non-Gaussian physics in the single-component Bose gas arises predominantly from three-body effects, Bose-Fermi mixtures already show non-Gaussian signatures at the more fundamental two-body level. Physically, this distinction arises from the fact that the fermions in the gas do not condense, meaning that the interspecies correlation functions can not be reduced to a Gaussian form by the emergence of a classical condensate field. Similarly, the interaction between a BEC and a filled Fermi sea generates pairs at nonzero total momentum, described by non-Gaussian correlation functions. This contrasts with the unitary Bose gas, where pair excitations from the condensate create zero-momentum pairs, which are naturally Gaussian. Due to the associated theoretical difficulties, previous studies of the dynamics of Bose-Fermi mixtures rely on various simplifications, such as large population imbalances or reductions of the dimensionality of the system Wouters et al. (2003); Bortolotti et al. (2005); Olsen et al. (2009); Cumby et al. (2013); Gärttner et al. (2019); Mistakidis et al. (2019); Abdullaev et al. (2019).
@@ II.1 Hamiltonian
where the primes denote annihilation of bosons and fermions in an energetically closed spin channel, and where $\phi(q)$ is the wave function of the closed-channel molecule as a function of the relative two-body momentum $q\equiv\absolutevalue{\mathbf{q}}$ |. The molecules have kinetic energy $\varepsilon_{\mathbf{Q}}^{M}=Q^{2}/2M$, and bare internal energy $\nu$ relative to the scattering threshold. A Bose-Fermi pair can associate to form a molecule, controlled by a separable interaction potential with coupling strength $g$ and form factor $\xi_{\mathbf{q}}=\Theta(\Lambda_{BF}-\absolutevalue{\mathbf{q}})$. Here $\Theta(x)$ is the Heaviside step function such that $\Lambda_{BF}$ defines an ultraviolet cutoff on the relative momentum, interpreted physically as the inverse of the range of the interspecies interaction.
@@ II.2 Cumulant expansion and equations of motion
Cumulant Notation Interpretation
$\expectationvalue*{\hat{b}_{\mathbf{0}}}_{\mathrm{c}}$ $\sqrt{V}\psi$ Condensate wave function.
$\expectationvalue*{\hat{b}_{\mathbf{k}}^{\dagger}\hat{b}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{B}$ Bosonic excitation density.
$\expectationvalue*{\hat{f}_{\mathbf{k}}^{\dagger}\hat{f}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{F}$ Fermionic density.
$\expectationvalue*{\hat{m}_{\mathbf{k}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $n_{\mathbf{k}}^{M}$ Molecular density.
$\expectationvalue*{\hat{b}_{\mathbf{k}}\hat{b}_{-\mathbf{k}}}_{\mathrm{c}}$ $\kappa_{\mathbf{k}}$ Bosonic pairing.
$\expectationvalue*{\hat{f}_{\mathbf{k}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $\chi_{\mathbf{k}}$ Atom-molecule conversion.
Closed-channel two-body
wave function.
$\expectationvalue*{\hat{b}_{\mathbf{k}-\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}^{\prime}}^{\dagger}\hat{m}_{\mathbf{k}}}_{\mathrm{c}}$ $M_{\mathbf{k},\mathbf{k}^{\prime}}$ Atom-molecule conversion.
$\expectationvalue*{\hat{b}_{\mathbf{k}-\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}^{\prime}}^{\dagger}\hat{f}_{\mathbf{k}}}_{\mathrm{c}}$ $L_{\mathbf{k},\mathbf{k}^{\prime}}$ Boson-fermion scattering.
Open-channel two-body
wave function.
Table 1: All cumulants included in our dynamical model, with a brief description of their physical interpretation. Note that by construction, a cumulant is only defined when all bosonic momenta are nonzero, except for the singlet $\expectationvalue*{\hat{b}_{\mathbf{0}}}_{\mathrm{c}}$.
@@
Next to the densities themselves, we also need to include the atom-molecule conversion processes that will act to source the density dynamics. The lowest order such process, shown in Fig. 1(a), describes the interaction between a fermion (solid black) and a boson in the BEC (wavy red), captured by the second order cumulant, or doublet, $\chi_{\mathbf{k}}$. To complete the set of doublet cumulants we define the bosonic pairing matrix $\kappa_{\mathbf{k}}$, which captures correlated pair excitations from the BEC. Note that this set excludes the fermionic and molecular pairing matrices $\expectationvalue*{\hat{f}_{\mathbf{k}}\hat{f}_{-\mathbf{k}}}_{\mathrm{c}}$ and $\expectationvalue*{\hat{m}_{\mathbf{k}}\hat{m}_{-\mathbf{k}}}_{\mathrm{c}}$, respectively, which always vanish provided the initial state for the fermionic component has a definite number of particles, and the Hamiltonian conserves the total fermionic population.
@@
In identical Bose and two-component Fermi gases, a truncation of the cumulant expansion at the level of second order correlations, or doublets, produces the Hartree-Fock-Bogoliubov approximation Blaizot and Ripka (1985), which is well studied in the literature and known to converge to the two-body Schrödinger equation in the short-range or vacuum limits Leggett (2006). In the Bose-Fermi mixture however, the same level of approximation leads to dynamical equations that do not correctly reproduce the vacuum two-body physics, pointing out the fundamental shortcomings of the doublet approximation and the inherent non-Gaussian nature of the pairing physics in the Bose-Fermi mixture Bortolotti et al. (2005, 2008). As we will show formally in Sec. III, this problem is remedied upon extending the doublet approximation with two non-Gaussian third order cumulants (triplets), defined in the last two rows of Table 1. Here, $L_{\mathbf{k},\mathbf{k}^{\prime}}$ captures the triplet contribution to the cumulant expansion of the total two-body correlator $\expectationvalue*{\hat{b}^{\dagger}\hat{f}^{\dagger}\hat{b}\hat{f}}$, alternatively interpreted as the open-channel analogue of the doublet $\chi_{\mathbf{k}}$. In turn, $M_{\mathbf{k},\mathbf{k}^{\prime}}$ captures the generalization of the scattering process in Fig. 1(a) to the case where the boson is in an excited state, as shown in Fig. 1(b). It is a necessary inclusion in the model as it ensures the conservation of the total system energy, or $\partialderivative{t}\expectationvalue*{\hat{H}}=0$.
@@
Figure 1: Two-body scattering diagrams that contribute to the density dynamics in the Bose-Fermi mixture, where red(black) wiggly lines represent (non)condensed bosons, and single(double) straight lines represent fermions(molecules). Process (a) represents scattering of fermions with the BEC, captured by $\chi_{\mathbf{Q}}$, while process (b) represents scattering of fermions with bosonic excitations, captured by $M_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}$. Here $\mathbf{Q}$ and $\mathbf{q}$ are the center-of-mass and relative momenta of the Bose-Fermi pair respectively.
@@
$\bar{n}_{\mathbf{k}}^{B}=\frac{1}{V}\sum_{\mathbf{k}^{\prime}}\zeta_{\mathbf{k}-\mathbf{k}^{\prime}}^{2}n_{\mathbf{k}^{\prime}}^{B},\qquad\bar{\kappa}=\frac{1}{V}\sum_{\mathbf{k}}\zeta_{2\mathbf{k}}\kappa_{\mathbf{k}}.$ (8)
@@
$i\dot{\psi}$ $=$ $v\left(\zeta_{0}^{2}\absolutevalue*{\psi}^{2}+2\bar{n}_{0}^{B}\right)\psi+v\psi^{*}\zeta_{0}\bar{\kappa}$ (9)
@@
Here the first line contains terms also found in the usual Gross-Pitaevskii equation for a single-component BEC with a contact potential of strength $v$, containing in addition the coupling to excited bosons and anomalous pairs through $n_{\mathbf{k}}^{B}$ and $\kappa_{\mathbf{k}}$ respectively Blaizot and Ripka (1985). The second line couples the BEC to the fermionic component, through the scattering process shown in Fig. 1(a). In turn, the dynamics of non-condensed bosons follows from,
@@
Here we have defined the Hartree-Fock Hamiltonian and pairing field Blaizot and Ripka (1985), quantifying the kinetic and interaction energies of bosonic pairs in the many-body background,
@@
$\Delta_{\mathbf{k}}$ $=$ $v\zeta_{2\mathbf{k}}\left[\zeta_{0}\psi^{2}+\bar{\kappa}\right].$ (13)
Note here that in the limit of vanishing Bose-Bose scattering length, $a_{BB}=0$, we have $h_{\mathbf{k}}^{\mathrm{hf}}=\varepsilon_{\mathbf{k}}^{B}$ and $\Delta_{\mathbf{k}}=0$. The coupling to the fermionic component is facilitated by the second line in Eq. (II.2), representing the conversion of an excited boson and a fermion into a molecule, as also shown diagramatically in Fig. 1(b). The analogous equations for the fermionic and molecular densities read,
@@
In both these equations the first and second line can be respectively connected to the diagrams in Figs. 1(a) and (b), capturing the interactions with the BEC and excited bosonic fractions. The cumulant $\chi_{\mathbf{k}}$ evolves as,
@@
$+$ $\frac{g}{\sqrt{V}}\sum_{\mathbf{k}^{\prime}}\xi_{\frac{m_{F}}{M}\mathbf{k}-\mathbf{k}^{\prime}}L_{\mathbf{k},\mathbf{k}^{\prime}}^{*}.$
As we will show in Sec. III, this equation can be most easily interpreted as the closed-channel component of the two-body Schrödinger equation for Bose-Fermi pairs in the gas, and is thus an essential inclusion in the model for obtaining the correct pairing physics.
@@
Here $M$ captures the atom-molecule conversion processes shown in Fig. 1b, as can be seen from the first term on the second line which takes the form of a Boltzmann type scattering term consistent with this diagram. The enhancement factors $(1+n_{\mathbf{k}-\mathbf{k}^{\prime}}^{B})$ and $(1-n_{\mathbf{k}^{\prime}}^{F}),(1-n_{\mathbf{k}}^{M})$ capture respectively the Bose enhancement and Pauli blocking in this process arising from the Bose and Fermi statistics Kira (2015). The scattering cumulant $L$ is essentially the open-channel analogue of $\chi$, and can thus be interpreted as the open-channel component of the two-body wave function for Bose-Fermi pairs, as we will show in Sec. III.
@@ III.1 Static pair energies
Before studying dynamics, we give a brief overview of relevant time-independent solutions to the embedded two-body Schrödinger equation. We will build up the complexity of the solution step by step, starting from an isolated Bose-Fermi pair interacting in vacuum, and subsequently adding in the many-body background. As we will show, it will be useful here to distinguish between stationary pairs whose total momentum $\mathbf{Q}=0$, and moving pairs with $\mathbf{Q}\neq 0$.
For an interacting Bose-Fermi pair in vacuum, the Hamiltonian [Eq. (1)] can be diagonalized analytically to obtain the binding energy $E_{\mathrm{vac}}$ of the dressed Feshbach dimer, resulting in the integral equation Kokkelmans et al. (2002),
@@
We now seek to extend Eq. (19) to a Bose-Fermi pair embedded in a BEC + Fermi sea mixture. In a many-body system, the Hamiltonian (1) is not analytically solveable in general. As pointed out in Ref. Dukelsky et al. (2011) however, the problem may be reduced to a solveable form by neglecting coupling terms between pairs that have nonzero center-of-mass momentum $\mathbf{Q}\neq 0$, equivalent to the replacement $\hat{m}_{\mathbf{Q}}\rightarrow\hat{m}_{0}\delta_{\mathbf{Q},0}$ in the Hamiltonian (see App. A). The exact eigenstates of this reduced many-body Hamiltonian that may be generated from our initial condition contain only a single interacting pair, formed from a boson in the BEC and a fermion at $\mathbf{k}=0$. The energy spectrum of this pair is obtained from the integral equation,
$E-\nu=g^{2}\xi_{\mathbf{0}}^{2}\frac{n_{0}^{B}}{E}+\frac{g^{2}}{V}\sum_{\mathbf{q}}\xi_{\mathbf{q}}^{2}\frac{1-n_{\mathbf{q}}^{F}}{E-\varepsilon_{\mathbf{q}}^{R}},$ (20)
which reduces to Eq. (19) in the vacuum limit $n_{B}\rightarrow 0$ and $n_{F}\rightarrow 0$, see App. A for more detail. The solutions $E$ of Eq. (20) are illustrated in Fig. 2, plotted as a function of the dimensionless ratio $(k_{F}a_{BF})^{-1}$ which measures the scattering length with respect to the Fermi momentum. Comparing with Eq. (19) we observe that the BEC and Fermi sea have competing effects on the energy of the pair. First, the Fermi sea blocks all momentum states $k<k_{F}$, thus raising the threshold for open-channel two-body scattering to $E_{\mathbf{0}}^{\mathrm{thr}}=k_{F}^{2}/2m_{R}$, a mechanism also responsible for Cooper pairing in metals Song et al. (2010); Song and Zhou (2011). The associated enhancement of the two-body interaction makes that the Feshbach dimer no longer disappears at resonance, but rather shifts above the vacuum threshold $E=0$ at positive scattering length $(k_{F}a_{BF})^{-1}=2/\pi$, as can be directly derived from Eq. (20) in the limit $n_{0}^{B}\rightarrow 0$. If the scattering length is tuned further to the attractive side of the resonance, the dimer moves up through the Fermi sphere towards the raised scattering threshold. Here we emphasize that this state constitutes an exact stationary state of the many-body Hamiltonian, meaning that the dimer state has infinite lifetime across the Feshbach resonance.